# Enhanced pedestal transport driven by edge collisionality on Alcator C-Mod and its role in regulating H-mode pedestal gradients

### v1

@@ Abstract
Experimental measurements of plasma and neutral profiles across the pedestal are used in conjunction with 2D edge modeling to examine pedestal stiffness in Alcator C-Mod H-mode plasmas. Experiments on Alcator C-Mod observed pedestal degradation and loss in confinement below a critical value of net power crossing the separatrix, $P_{\mathrm{net}}=P_{\mathrm{net}}^{\mathrm{crit}}\approx 2.3$ MW. New analysis of ionization and particle flux profiles reveal saturation of the pedestal electron density, $n_{e}^{\mathrm{ped}}$ despite continuous increases in ionization throughout the pedestal, inversely related to $P_{\mathrm{net}}$. A limit to the pedestal $\nabla n_{e}$ emerges as the particle flux, $\Gamma_{D}$ continues to grow, implying increases in the effective particle diffusivity, $D_{\mathrm{eff}}$. This is well-correlated with the separatrix collisionality, $\nu^{*}_{\mathrm{sep}}$ and a turbulence control parameter, $\alpha_{t}$, implying a possible transition in type of turbulence. The transition is well correlated with the experimentally observed value of $P_{\mathrm{net}}^{\mathrm{crit}}$. SOLPS-ITER modeling is performed for select discharges from the power scan, constrained with experimental electron and neutral densities, measured at the outer midplane. The modeling confirms general growth in $D_{\mathrm{eff}}$, consistent with experimental findings, and additionally suggests even larger growth in $\chi_{e}$ at the same $P_{\mathrm{net}}^{\mathrm{crit}}$.
@@ 1 Introduction
The physics that determine plasma profiles in the edge of tokamak plasmas remains among the most elusive in tokamak physics. It is known that at sufficient input power, a transition to a high-confinement mode (H-mode) occurs, resulting in a substantial reduction in edge transport, increasing pressure gradients and forming the so-called “pedestal” [1]. It is well-known that achievable core pressures are very sensitive to the conditions at the boundary [2, 3, 4, 5, 6]. Next-generation devices [7, 8] seeking to maximize fusion gain rely on a robust edge pedestal. If the H-mode pedestal is to be exploited for high fusion gain, confidence in understanding the mechanisms determining its structure and the ability to form a desired type of pedestal must increase, especially in view of integrating this edge with a viable heat exhaust SOL and divertor solution.
Some models have had reasonable success in predicting the pressure pedestal structure of H-modes exhibiting Type-I edge localized modes (ELMs) in conventional aspect ratio machines [9, 10, 11, 12]. Regardless, prediction of pedestal structure in turbulence-limited H-modes without ELMs is less certain [13]. Density pedestals in particular are especially poorly understood, largely as a result of a lack of edge ionization source measurements. It is unclear to what extent neutral fueling, as opposed to particle transport, is responsible for the build up of edge pedestal density gradients [14, 15]. Certain models suggest that the density pedestal width may be strongly linked to the neutral penetration depth [16, 17]. While this may hold in some H-mode regimes, it cannot explain the width of all density pedestals [14, 18]. In ohmic L-modes, turbulent transport has been identified as important for regulating edge gradients [19, 20], and it is unknown to what degree turbulence is responsible at holding H-mode gradients at or near marginal stability. Since the very small length scales in the pedestal makes direct gyrokinetic simulation of these instabilities challenging, reduced-order models can serve as alternatives. Knowledge of the distribution of sources in the edge, in particular the ionization source, is an asset to testing reduced-order fueling and transport models.
@@
Alcator C-Mod operated until 2016, and it routinely achieved the highest plasma densities of current tokamaks, fairly close to and often exceeding densities proposed for reactors [25, 26, 15]. As edge plasma density rises, opaqueness to neutrals increases, and fueling through gas puffing may become a challenge. As opaqueness increases, the ionization profile may move radially outward onto the unconfined plasma. If this is the case for future reactors, their ability to sustain large density gradients in the edge may be significantly reduced. A proposed metric for plasma opaqueness, $\eta=a\times n_{e}$, where $a$ is the plasma minor radius and $n_{e}$ is the electron density, puts Alcator C-Mod’s opaqueness close to that expected on SPARC and ITER [15, 27]. Higher $a$ and $n_{e}$ are thought to increase the difficulty with which an edge neutral can reach the core, increasing the plasma’s opaqueness. Ionization profiles on Alcator C-Mod may thus provide insight into what ionization may look like in future devices. Furthermore, pedestal resiliency to fueling and evidence of critical edge gradients [18, 19] on C-Mod may be a feature of high-density devices with which future reactors need to contend.
@@
This paper begins by outlining the set of diagnostics used to create a database of edge ionization and pedestal profiles. It then compares experimental ionization measurements with density measurements from a particular set of experiments at fixed field, current, and plasma shape to make comments about observed pedestal stiffness, and the accompanying increase in particle transport. Section 4 then considers the pedestal operational space, and makes claims about how the C-Mod plasma moves around this operational space. It identifies collisionality, particularly at the edge, as a key driver for determining both the density and temperature pedestal. To supplement these experimental observations, the next section focuses on results from a set of simulations, containing high-fidelity and experimentally-validated 2D calculations of the ionization source. The paper then ends with discussion on implications of these results for next-generation devices, and future steps that will be taken to develop a predictive model for next-generation density pedestals.
@@ 2 Experimental ionization and particle flux profiles in a set of magnetic balance experiments
In 2007, a series of experiments were performed on Alcator C-Mod to explore techniques for decoupling particle transport from energy transport in H-mode pedestals. One particular experiment sought to modify magnetic balance to study its effect on the enhanced Dα (EDA) H-mode regime. The EDA H-mode is a type of H-mode not limited by peeling-ballooning MHD modes, as the more typical type-I ELMing H-modes are [28, 21]. To study the effect of the magnetic drift on EDA H-mode pedestal quality, for a fixed toroidal magnetic field, $B_{t}$ = 5.4 T, plasma current, $I_{P}$ = 0.8 MA, and $\nabla B$-drift direction, the active null was moved from the lower chamber to the upper chamber, effectively modifying the outer midplane distance between the separatrices passing through the two X-points, $\Delta R_{sep}$, from around -5 mm to +5 mm. The current work, however, focuses solely on plasmas solidly in single null, with $\Delta R_{sep}\approx-5$ mm. These were all obtained at relatively fixed shape, with elongation, $\kappa\approx 1.6$, and triangularity, $\delta\approx 0.5$. They had a major radius, $R_{0}\approx 0.67$ m and a minor radius, $a\approx 0.22$ m. Figure 1 shows a typical magnetic equilibrium for these plasmas, showing also views of the diagnostics described in to be described in Sections 2.2 and 2.3.
This paper presents an integrated analysis of particle and neutral pedestal transport from these experiments. It does so by combining radially-resolved edge electron density and temperature ($n_{e}$, $T_{e}$) measurements from Thomson Scattering (TS) with neutral inferences of neutral density ($n_{0}$), ionization source ($S_{\mathrm{ion}}$), and radial particle flux ($\Gamma_{D}$) from a Lyman-$\alpha$ (Lyα) camera.
Figure 1: Typical magnetic equilibrium for LSN discharges in effective power scan. The thick red line demarcates the last closed flux surface, separating the closed field lines, shown in cyan, and the open field lines, shown in dark blue. The TS scattering system views vertically down the plasma. The sparser core system (CTS) views across the majority of the core plasma. ETS views the area around the separatrix, located poloidally near the crown of the plasma in the LSN configuration. The LYMID system, shown in bright blue, is located exactly at the midplane and views toroidally along 20 chords.
@@ 2.1 Pedestal degradation with net power
Figure 2 shows the large changes that occur in these discharges as $P_{\mathrm{net}}\approx 2.3$ MW. Figures 2a - d show that the line-integrated density, $\overline{n}_{e}$, drops, concurrent with large increases in $W_{\mathrm{MHD}}$, the maximum normalized pressure gradient, $\alpha_{\mathrm{MHD}}$, and the normalized confinement factor, $H_{98}$. Figures 2e - g show changes in pedestal parameters as measured by the edge Thomson Scattering (ETS) diagnostic, outlined in Section 2.3. There is a large drop in the pedestal collisionality, $\nu^{*}_{95}$, resulting from a very slight drop in $n_{e}^{\mathrm{ped}}$, but a large rise in $T_{e}^{\mathrm{ped}}$. Revisiting this particular experiment through inspection of edge ionization profiles reveals the role of plasma transport in regulating the pedestal and thus, the H-mode quality. Figures 2h - k allude to this, showing large drops in the inferred neutral density, $n_{0}$, ionization source, $S_{\mathrm{ion}}$, particle flux, $\Gamma_{D}$, and particle diffusivity, $D_{\mathrm{eff}}$, all at the top of the pedestal. The discharges used in this dataset are listed below  [^1] Shot numbers in dataset: 1070821003, 1070821004, 1070821005, 1070821008, 1070821009, 1070821012, 1070821013, 1070821020, 1070821021, 1070821022, 1070821023, 1070821025, and 1070821029.
@@ 2.2 Neutral emissivity from Lyα camera
Radial profiles of $n_{0}$, $S_{\mathrm{ion}}$, and $\Gamma_{D}$ are constructed from a midplane Lyα camera called LYMID. It was installed in 2007 as the successor to an earlier Lyα system [29, 30]. It was located just below the midplane ($Z=-0.04$ m) and viewed toroidally. Its 20 chords were tangent to flux surfaces covering a region in the radial direction of just over 5 cm at the plasma edge. The LYMID diagnostic collected Lyα light from the 121.6 nm visible ultra-violet light transition from the edge population of neutral deuterium atoms. This LYMID diagnostic recorded line-integrated brightness measurements from each of its viewing chords as a function of their tangency radius. We use the same tomographic inversion algorithm developed for LLAMA, a similar system on DIII-D [31, 32] to invert the brightness profile. This yields the Lyα emissivity, $\epsilon_{\mathrm{Ly}_{\alpha}}$, as a function simply of radius, and no longer of toroidal angle, under the assumption of axisymmetric emission.
Figure 2: Various experimentally-determined quantities plotted against $P_{\mathrm{net}}$. (a) - (d) show quantities typically linked with plasma confinement. (e) - (g) show variation of plasma quantities near the pedestal top, and (h) - (k) show variation of quantities more directly related to neutrals. There are clear changes in all quantities at a critical value of $P_{\mathrm{net}}$.
@@ 2.3 Edge Thomson scattering for plasma profile analysis
In addition to neutral emissivity measurements, we use the edge Thomson scattering (ETS) diagnostic, which was routinely used to diagnose the edge plasma and in particular, the pedestal on C-Mod. ETS measured both electron density, $n_{e}$ and electron temperature, $T_{e}$ about a region spanning 3 cm of the plasma edge. Measurements were located in the upper chamber near the crown of lower single null (LSN) plasmas, and could diagnose $n_{e}$ and $T_{e}$ profiles with order millimeter resolution when mapped to the midplane [33]. In order to estimate gradients as well as to facilitate interpolation onto the radial coordinate basis of the neutral emissivity profile, a fit to both $n_{e}$ and $T_{e}$ is applied. It was determined that a hyperbolic tangent as in [34], but supplemented with polynomial terms to allow for flexibility in adjusting the gradient inside the pedestal and in the SOL. This provided reasonably good fits to the plasma data. Equation 1 shows the functional form of the modified hyperbolic tangent (mtanh).
@@
Since the mtanh has a closed form, the $n_{e}$ and $T_{e}$ gradients ($\nabla n_{e}$ and $\nabla T_{e}$) can be computed analytically using the fit coefficients. A bootstrapping technique is then used to compute uncertainties in the fitted profiles. This is done by perturbing the experimental ETS data points, refitting, and computing statistics on the fits. The perturbation to the experimental points is sampled from a Gaussian distribution with a half-width equal to the experimental uncertainty of the ETS data point. The fitting process is repeated multiple times and the uncertainty is set to the standard deviation of the ensemble of perturbed fits. This technique enables uncertainty estimation not only in the absolute $n_{e}$ and $T_{e}$ profiles, but also in their gradients. It is described in detail in [35].

### v2

@@ Abstract
Experimental measurements of plasma and neutral profiles across the pedestal are used in conjunction with 2D edge modeling to examine pedestal stiffness in Alcator C-Mod H-mode plasmas. Enhanced Dα (EDA) experiments on Alcator C-Mod observed pedestal degradation and loss in confinement below a critical value of net power crossing the separatrix, $P_{\mathrm{net}}=P_{\mathrm{net}}^{\mathrm{crit}}\approx 2.3$ MW, in the absence of any external fueling. New analysis of ionization and particle flux profiles reveal saturation of the pedestal electron density, $n_{e}^{\mathrm{ped}}$, despite continuous increases in ionization throughout the pedestal, inversely related to $P_{\mathrm{net}}$. A limit to the pedestal $\nabla n_{e}$ emerges as the particle flux, $\Gamma_{D}$, continues to grow, implying increases in the effective particle diffusivity, $D_{\mathrm{eff}}$. This is well-correlated with the separatrix collisionality, $\nu^{*}_{\mathrm{sep}}$ and a turbulence control parameter, $\alpha_{t}$, implying a possible transition in type of turbulence. The transition is well correlated with the experimentally observed value of $P_{\mathrm{net}}^{\mathrm{crit}}$. SOLPS-ITER modeling is performed for select discharges from the power scan, constrained with experimental electron and neutral densities, measured at the outer midplane. The modeling confirms general growth in $D_{\mathrm{eff}}$, consistent with experimental findings, and additionally suggests even larger growth in $\chi_{e}$ at the same $P_{\mathrm{net}}^{\mathrm{crit}}$.
@@ 1 Introduction
The physics that determine plasma profiles in the edge of tokamak plasmas remains among the most elusive in tokamak physics. It is known that at sufficient input power, a transition to a high-confinement mode (H-mode) occurs, resulting in a substantial reduction in edge transport, increasing pressure gradients, and forming the so-called “pedestal" [1]. It is well-known that achievable core pressures are very sensitive to the conditions at the boundary [2, 3, 4, 5, 6]. Next-generation devices [7, 8] seeking to maximize fusion gain rely on a robust edge pedestal. If the H-mode pedestal is to be exploited for high fusion gain, confidence in understanding the mechanisms determining its structure and the ability to form a desired type of pedestal must increase, especially in view of integrating this edge with a viable heat exhaust scrape-off layer (SOL) and divertor solution.
Some models have had reasonable success in predicting the pressure pedestal structure of H-modes exhibiting Type-I edge localized modes (ELMs) in conventional aspect ratio machines [9, 10, 11, 12]. Regardless, prediction of pedestal structure in turbulence-limited H-modes without ELMs is less certain [13]. Density pedestals in particular are especially poorly understood, largely as a result of a lack of edge ionization source measurements. It is unclear to what extent neutral fueling, as opposed to particle transport, is responsible for the build up of edge pedestal density gradients [14, 15]. Certain models suggest that the density pedestal width may be strongly linked to the neutral penetration depth [16, 17]. While this may hold in some H-mode regimes, it cannot explain the width of all density pedestals [14, 18]. In ohmic L-modes, turbulent transport has been identified as important for regulating edge gradients [19, 20], and it is unknown to what degree turbulence is responsible for holding H-mode gradients at or near marginal stability. Since the very small length scales in the pedestal makes direct gyrokinetic simulation of these instabilities challenging, reduced-order models can serve as alternatives. Knowledge of the distribution of sources in the edge, in particular the ionization source, is an asset to testing reduced-order fueling and transport models.
@@
Alcator C-Mod operated until 2016, and it routinely achieved the highest plasma densities of current tokamaks, fairly close to and often exceeding densities proposed for reactors [25, 26, 15]. As edge plasma density rises, opaqueness to neutrals increases, and fueling through gas puffing may become a challenge. As opaqueness increases, the ionization profile may move radially outward onto the unconfined plasma. If this is the case for future reactors, their ability to sustain large density gradients in the edge may be significantly reduced. A proposed metric for plasma opaqueness, $\eta=a\times n$, where $a$ is the plasma minor radius and $n$ is the plasma density, puts Alcator C-Mod’s opaqueness close to that expected on SPARC and ITER [15, 27]. Higher $a$ and $\overline{n}_{e}$ are thought to increase the difficulty with which an edge neutral can reach the core. Ionization profiles on Alcator C-Mod thus provide insight into how ionization could look in future devices. Furthermore, pedestal resiliency to fueling and evidence of critical edge gradients [18, 19] on C-Mod may be a feature of high-density devices with which future reactors need to contend.
@@
This paper begins by outlining the set of diagnostics used to create a database of edge ionization and pedestal profiles. Section 3 then compares experimental ionization measurements with density measurements from a particular set of experiments at fixed magnetic field, plasma current, and plasma shape to make comments about observed pedestal stiffness, and the accompanying increase in particle transport. Section 4 then considers the pedestal operational space, and makes claims about how the C-Mod plasma moves around this operational space. It identifies collisionality, particularly at the edge, as a key driver for determining both the density and temperature pedestal. To supplement these experimental observations, Section 5 focuses on results from a set of simulations containing high-fidelity and experimentally-validated 2D calculations of the ionization source. The paper then ends with discussion on implications of these results for next-generation devices, and future steps that will be taken to develop a predictive model for next-generation density pedestals.
@@ 2 Experimental ionization and particle flux profiles in a set of magnetic balance experiments
In 2007, a series of experiments were performed on Alcator C-Mod to study the effect of switching the magnetic geometry from lower single null (LSN) to upper single null (USN) on the EDA H-mode regime. The EDA H-mode is a type of H-mode not limited by peeling-ballooning MHD modes, as the more typical type-I ELMing H-modes are [28, 21]. To study the effect of the magnetic drift on EDA H-mode pedestal quality, for a fixed toroidal magnetic field, $B_{t}$ = 5.4 T, plasma current, $I_{P}$ = 0.8 MA, and $\nabla B$-drift direction, the active null was moved from the lower chamber to the upper chamber, effectively modifying the outer midplane distance between the separatrices passing through the two X-points, $\Delta R_{\mathrm{sep}}$, from around -5 mm to +5 mm. The current work focuses solely on plasmas solidly in LSN, with $\Delta R_{\mathrm{sep}}\approx-5$ mm. These were all obtained at relatively fixed shape, with elongation, $\kappa\approx 1.6$, and triangularity, $\delta\approx 0.5$. They had a major radius, $R_{0}\approx 0.67$ m and a minor radius, $a\approx 0.22$ m. Figure 1 shows a typical magnetic equilibrium for these plasmas, showing also views of the diagnostics to be described in Sections 2.2 and 2.3. As was the case for many H-modes on Alcator C-Mod, gas was only injected in the L-mode phase, and the target density was sustained in the H-mode without any additional gas request on the controller. These plasmas also received no external impurity injection.
This paper presents a self-consistent analysis of particle and neutral pedestal transport from these experiments. It does so by combining radially-resolved edge electron density and temperature ($n_{e}$, $T_{e}$) measurements from Thomson Scattering (TS) with measurements of neutral emissivity from a Lyman-$\alpha$ camera, $\epsilon_{\mathrm{Ly_{\alpha}}}$, to infer neutral density, $n_{0}$, ionization source, $S_{\mathrm{ion}}$, radial particle flux, $\Gamma_{D}$, and effective diffusivity, $D_{\mathrm{eff}}$.
Figure 1: Typical magnetic equilibrium for LSN discharges in effective power scan. The thick red line demarcates the last closed flux surface, separating the closed field lines, shown in cyan, and the open field lines, shown in dark blue. The TS system in purple views vertically down the plasma. The sparser core system views across the majority of the core plasma. The edge TS system (ETS) views the area around the separatrix, located poloidally near the crown of the plasma in the LSN configuration. The LYMID system, shown in bright blue, is located exactly at the midplane and views toroidally along 20 chords.
@@ 2.1 Pedestal degradation with net power
Figure 2 shows an effective scan in $P_{\mathrm{net}}$ from 1.3 MW to 3.4 MW. It shows large changes to many quantities in these discharges as $P_{\mathrm{net}}$ approaches and crosses a critical value of $P_{\mathrm{net}}$, $P_{\mathrm{net}}^{\mathrm{crit}}\approx$ 2.3 MW. The left panel of Figure 2 shows that the line-integrated density, $\overline{n}_{e}$, drops by $\sim$8%, simultaneous with large increases in $W_{\mathrm{MHD}}$, the maximum normalized pressure gradient, $\alpha_{\mathrm{MHD}}$, and the normalized confinement factor, $H_{\mathrm{98,y2}}$ ($H_{98}$) [29]. Recalling the expression for the opaqueness heuristic introduced in Section 1, the plot of $\overline{n}_{e}$ also gives information about the opaqueness in these shots. Multiplying $\overline{n}_{e}$ by $a$ for these shots gives a range in opaqueness from $5.5-6.8\times 10^{19}$m-2. While these values are among the highest of current devices, they are not the highest achieved on C-Mod. For example, plasmas in the final campaigns of C-Mod achieved opaqueness values of up to $1.1\times 10^{20}$m-2 [30]. For reference, recent plasmas with Greenwald fraction larger than unity on DIII-D achieved an opaqueness just under $4.4\times 10^{19}$m-2 [31]. ITER and SPARC, on the other hand, are expected to operate at opaqueness values of $2.2\times 10^{20}$ m-2 and $1.7\times 10^{20}$ m-2, respectively [32, 6]. If instead, the plasma density is taken as $n=\frac{1}{2}(n_{e}^{\mathrm{sep}}+n_{e}^{\mathrm{ped}})$, as suggested in [15], the opaqueness for these discharges ranges from from $3.1-4.3\times 10^{19}$m-2 and that predicted for SPARC and ITER are $1.6\times 10^{20}$ m-2 and $1.5\times 10^{20}$ m-2, respectively [33, 34].
The center panel of Figure 2 shows changes in pedestal plasma parameters as measured by the ETS diagnostic, outlined in Section 2.3. There is a large drop in the collisionality at $\psi_{n}$ = 0.95, $\nu^{*}_{95}$, resulting from a very slight drop in $n_{e}^{\mathrm{ped}}$, but a large rise in $T_{e}^{\mathrm{ped}}$. Each of these pedestal quantities is computed at their respective pedestal top, defined by the fit function described in Section 2.3 rather than at a particular flux surface. In general, these are at different locations for $n_{e}$ and $T_{e}$. The bottom plot of the center panel differentiates, with color, discharges that are at high values of $\nu^{*}_{95}>2$ and low values of $\nu^{*}_{95}<2$. This transition in collisionality occurs at around a value of $P_{\mathrm{net}}\approx$ 2.3 MW, at which point all the quantities in Figure 2 show clear changes.
Revisiting this particular experiment through inspection of edge ionization profiles reveals the role of plasma transport in regulating the pedestal and thus, the H-mode quality. The right panel of Figure 2 alludes to this, showing large drops in the inferred $n_{0}$, $S_{\mathrm{ion}}$, $\Gamma_{D}$, and $D_{\mathrm{eff}}$, all at the top of the pedestal. The discharges used in this dataset are the following: 10708210[03/04/05/08/09/12/13/20/21/22/23/25/29].
@@ 2.2 Neutral emissivity from Lyα camera
Radial profiles of $n_{0}$, $S_{\mathrm{ion}}$, $\Gamma_{D}$, and $D_{\mathrm{eff}}$ are reconstructed from a midplane Lyα camera called LYMID, using the methodology outlined in Section 2.4. The camera was installed in 2007 as the successor to an earlier Lyα system [35, 36]. It was located just below the midplane ($Z=-0.04$ m) and consisted of toroidal views. Its 20 chords were tangent to flux surfaces covering a region in the radial direction of just over 5 cm at the plasma edge. The LYMID diagnostic collected Lyα light from the 121.6 nm vacuum ultra-violet light transition from the edge population of neutral deuterium atoms. This LYMID diagnostic recorded line-integrated brightness measurements from each of its viewing chords as a function of their tangency radius. We use the same tomographic inversion algorithm developed for LLAMA, a similar system on DIII-D [37, 38] to invert the brightness profile. This yields $\epsilon_{\mathrm{Ly}_{\alpha}}$ as a function simply of radius, and no longer of toroidal angle, under the assumption of axisymmetric emission. No radial shift is applied to $\epsilon_{\mathrm{Ly}_{\alpha}}$. Instead, an error associated with the magnetic reconstruction is included, as delineated at the end of Section 2.4.
Figure 2: Various experimentally-determined quantities plotted against $P_{\mathrm{net}}$. The plots in the left panel show quantities typically linked with plasma confinement in dark blue circles. The center panels show variation of plasma quantities at the pedestal top in gray triangles, with the exception of the bottom panel, showing collisionality at $\psi_{n}=0.95$, for which a distinction in color is made. Purple triangles have $\nu^{*}_{95}>2$, while green triangles have $\nu^{*}_{95}<2$. This transition value roughly coincides with a critical value of $P_{\mathrm{net}}\approx$ 2.3 MW, at which all other quantities exhibit step-like changes. Finally, the panel on the right shows variation in quantities directly related to neutrals using light blue squares.
@@ 2.3 Edge Thomson scattering for plasma profile analysis
In addition to neutral emissivity measurements, we use the ETS diagnostic, which was routinely used to diagnose the edge plasma, and in particular, the pedestal, on C-Mod. ETS measured both $n_{e}$ and $T_{e}$ in a region spanning 3 cm of the plasma edge. Measurements were located in the upper chamber near the crown of LSN plasmas, and could diagnose $n_{e}$ and $T_{e}$ profiles with order millimeter resolution when mapped to the midplane [39]. In order to estimate gradients as well as to facilitate interpolation onto the radial coordinate basis of the neutral emissivity profile, a fit to both $n_{e}$ and $T_{e}$ is applied. It was determined that a hyperbolic tangent as in [40], but supplemented with polynomial terms to allow for flexibility in adjusting the gradient inside the pedestal and in the SOL, was appropriate to fit the plasma data. Equation 1 shows the functional form of the modified hyperbolic tangent (mtanh).
@@
For ETS, we use the two-point model [41] to align $n_{e}$ and $T_{e}$ relative to the separatrix, $\psi_{n}=1$. This model assumes that the power carried by the electrons in the SOL leaves the core entirely at the outer midplane and is transported to the divertor target via conduction only. It uses the Spitzer-Härm conductivity, which is proportional to the parallel temperature gradient in the SOL. Hence, the equality,
can be integrated to find $T_{e}^{\mathrm{sep}}$ at the outer midplane (OMP) [41]. In this equation, $\kappa_{0,e}$ is the Spitzer-Härm conductivity, $R$ is the plasma major radius, $\lambda_{q}$ is the width of the heat flux channel, $B_{\theta}$ is the poloidal magnetic field. In the absence of measurements of the parallel heat flux width, we use the Brunner scaling for $\lambda_{q}$ [42]. We further assume negligible temperature at the divertor (i.e. $T_{e,\mathrm{target}}^{7/2}\ll T_{e,\mathrm{OMP}}^{7/2}$), as some target data available for these discharges indicate that $T_{e,\mathrm{target}}$ does not typically exceed 25 eV for even the highest values of $P_{\mathrm{net}}$. We also estimate the connection length required for the integral as $L_{\parallel}=q_{95}\pi R$, where $q_{95}$ is the safety factor at the 95% flux surface. Finally, the factor of $\frac{1}{2}$ in front of $P_{\mathrm{net}}$ represents the assumption that half of the power crossing the separatrix is carried by the electrons (and half by the ions), an assumption that will be discussed further in the modeling setup in Section 5. With $T_{e,\mathrm{OMP}}^{\mathrm{sep}}$ in hand, we find the radial shift required to align the fit to the $T_{e}$ profile at $\psi_{n}=1$ and apply the same shift to $n_{e}$ before using these shifted kinetic profiles to calculate the rate coefficients used for the neutral inferences.